The measure of the second triangle is 21 degrees more than the measure of the first angle. The measure of the third angle is 15 degrees less than the measure of the first angle. Find the meausre of each angle of the triangle.
Answer by rfer(12730)

x+x+21+x-15=180
3x=174
x=58 degrees
x+21=79 degrees
x-15=43 degrees